Find a solution to the system of equations by graphing.
2x-y=8
x+y=1

Answer by mananth(16946)

2x-y=8....................I
x+y=1.......................II

Find the x & y intercept points by plugging x=0,y=0 in both equations.
equation I
when x= 0 y= -8 ,( 0 , -8 )
when y = 0 x= 8 ,( 4 , 0 )
Equation 2
when x= 0 y= 1 ( 0 , 1 )
when y = 0 x= 1 ( 1 , 0 )

The solution is(3,-2)